# A Purely Magnetic Route to High-Harmonic Spin Pumping

### v1

@@ Abstract
Spin pumping provides a fundamental route for dynamical spin transport, yet in its conventional form it produces only linear spin responses at the driving frequency. Recent studies have shown that spin-orbit coupling (SOC) can lift these restrictions and enable highly nonlinear spin and charge currents. Here we propose a distinct mechanism for high-harmonic spin pumping that does not rely on spin–orbit interactions. We show that nonlinearities in the adiabatic energy spectrum—rather than SOC itself—constitute the essential ingredient for high-harmonic generation in pumped spin currents. Such nonlinearities can arise in purely magnetic systems when a secondary magnetic order parameter is introduced perpendicular to the cone axis of a precessing magnetic moment. As a result, spin pumping and its higher harmonics emerge even in the complete absence of SOC. Our findings establish a new route to ultrafast spin pumping based solely on magnetic structure and dynamics.
@@
Spin pumping has been a cornerstone of spintronics since its formulation in the early 2000s [1, 2]. It provides a mechanism by which a time-dependent magnetization injects a spin current into an adjacent conductor—most commonly a normal metal—leading to enhanced Gilbert damping and enabling electrical detection via spin–charge conversion mechanisms where spin currents can be detected as transverse charge voltages [3, 4, 5, 6]. More recently, both theoretical [7] and experimental [8] studies have pointed to the possibility of longitudinal spin pumping, indicating that dynamical spin transport can extend beyond the conventional transverse charge recording paradigm.
@@
In parallel, it has been shown that magnetization dynamics can induce strongly nonlinear transport responses, including the generation of ultrahigh harmonics in spin and charge pumping when additional symmetry-breaking mechanisms are present [9, 10]. Such high-harmonic spin-current emission naturally connects spin pumping to the broader field of ultrafast and nonlinear spintronics, where spin currents driven on femtosecond to picosecond timescales enable efficient terahertz radiation and frequency conversion [11, 12, 13, 14]. From this perspective, spin pumping emerges as a platform at the intersection of high-harmonic generation (HHG)[15] and spintronic terahertz emission [16], underscoring the intrinsically multidisciplinary nature of dynamical carrier transport phenomena.
@@
In its conventional realization, spin pumping is driven by ferromagnetic [2] or antiferromagnetic [17] resonance, and the emitted spin current oscillates at the same frequency as the driving dynamics. Symmetry considerations prohibit direct charge transport in this setting, such that only pure spin angular momentum is pumped. The effect has been explored across a wide range of material platforms and geometries, including bilayers and low-dimensional systems, with growing interest in its role for precision spin-transport measurements and high-frequency spin-current generation [18, 19, 20]. In a recent proposal [9], we have demonstrated that extreamly high harmonics can emerge in carrier pumping in the presence of SOC of either relativistic or nonrelativistic nature [21]. In the present letter, we propose a conceptually distinct route to high-harmonic spin pumping that does not rely on SOC. Building on recent insights into the role of nonlinearities in adiabatic energy spectra [10], we show that the essential ingredient underlying high-harmonic pumping is not SOC per se, but rather the presence of intrinsic nonlinear band dynamics. While such nonlinearities often arise in systems with spin–orbit interactions, we demonstrate that they can also be generated in purely magnetic systems. Specifically, we consider a SOC-free magnetic structure in which a second static order parameter is introduced perpendicular to the cone axis of the precessing magnetization. This additional magnetic order qualitatively reshapes the adiabatic energy levels, inducing strong nonlinear features that directly propagate into the dynamical transport response.
As a result, spin pumping accompanied by a cascade of higher harmonics becomes possible even in the complete absence of SOC. Our findings establish a new mechanism for nonlinear and high-frequency spin pumping rooted solely in magnetic structure and dynamics, thereby broadening the class of systems capable of hosting ultrafast spin-transport phenomena beyond the existing spin–orbit-based approaches.
I From Band Dynamics to Nonlinear Spin Transport
@@
In conventional spin pumping, a precessing magnetic order placed in proximity to a normal metal injects a spin current into the latter. From a band-structure perspective, this process is governed by the adiabatic evolution of electronic states in a time-dependent magnetic background. To identify the conditions under which this evolution can generate high-harmonic responses in transport, it is sufficient to consider a minimal model consisting of a one-dimensional chain hosting a driven magnetic order. Despite its simplicity, this model provides a transparent platform to connect nonlinear band dynamics to nonlinear spin transport.
@@
II Numerical Simulations
@@
To complement the analytical analysis of the energy spectrum presented above, we now turn to numerical simulations of the time-dependent transport response generated by the Hamiltonian in Eq. (3). In particular, we consider a one-dimensional chain hosting a driven magnetic order and compute the resulting spin and charge currents flowing into an attached normal-metal lead. This approach allows us to directly access the full time-resolved transport response beyond analytical limits and to validate the emergence of nonlinear pumping signatures discussed previously.
@@
[…]
@@
$\boldsymbol{\rm{I}}^{\sigma}(t)=A_{r}\,\boldsymbol{m}(t)\times\dot{\boldsymbol{m}}(t)-A_{i}\,\dot{\boldsymbol{m}}(t),$ (6)
@@
We now turn to the case where the static transverse magnetic field is included. Figure 2 shows the time-domain signals and corresponding Fourier spectra for two different precession cone angles. In agreement with spin-pumping theory, the dc charge current remains zero [see Figs. 2(b) and 2(d)]. However, the presence of the transverse field gives rise to a strongly nonlinear dynamical response, as evidenced by the emergence of higher harmonics in the Fourier spectra shown in Figs. 2(a) and 2(c).
It is also worth noting that the spin current polarized along the direction of the static field exhibits a vanishing dc component. In contrast, the spin currents polarized along the $y$ and $z$ directions acquire finite dc contributions. For a small precession cone angle, $\theta=\pi/8$, only a limited number of harmonics is observed. When the dynamical magnetization maximally couples to the secondary magnetic order, corresponding to $\theta=\pi/2$, the harmonic spectrum extends up to the 30th order [Fig. 2(c)].
@@
Figure 3: Calculated dc components of the pumped spin current polarized along the $y$ [(a)] and $z$ [(b)] directions are shown as functions of the driving frequency $\omega$ for precession cone angles ranging from $\theta=\pi/8$ to $\pi/2$. The pumped currents are evaluated for $J_{0}=J_{1}=2\gamma$. The dashed brown and black lines correspond to linear fits of the underlying data.
III Discussion and Conclusion
@@
Another notable feature of this mechanism is the emergence of dc spin components in selected polarization channels, while the dc charge current remains forbidden by symmetry. This behavior clearly distinguishes the present nonlinear magnetic pumping regime from conventional adiabatic spin pumping and provides an experimentally accessible signature of the underlying nonlinear dynamics.
Taken together, these results establish a conceptually new route to high-harmonic spin pumping based solely on magnetic nonlinearities. The simplicity of the model and the robustness of the effect suggest that such nonlinear transport phenomena should be realizable in a broad class of magnetic systems, providing a promising platform for ultrafast spintronic functionalities and high-frequency signal generation.

### v2

@@ Abstract
Spin pumping provides a fundamental route for dynamical spin transport, yet in its conventional form it produces only linear spin responses at the driving frequency. Recent studies have shown that spin–orbit coupling (SOC) can lift these restrictions and enable strongly nonlinear spin and charge currents. Here we propose a distinct mechanism for high-harmonic spin pumping that operates independently of spin–orbit interactions. We demonstrate that purely magnetic structures can sustain high-harmonic spin currents when an additional static magnetic order parameter is introduced perpendicular to the cone axis of a precessing magnetization. This secondary magnetic order qualitatively reshapes the dynamical response, giving rise to a cascade of higher harmonics in the pumped spin current in the absence of SOC. Our results establish a SOC-independent route to ultrafast and nonlinear spin pumping rooted solely in magnetic structure and dynamics.
I Introduction
@@
Spin pumping has been a cornerstone of spintronics since its formulation in the early 2000s [1, 2]. It provides a mechanism by which a time-dependent magnetization injects a spin current into an adjacent conductor—most commonly a normal metal—leading to enhanced Gilbert damping and enabling electrical detection via spin–charge conversion mechanisms where spin currents can be detected as transverse charge voltages [3, 4, 5, 6]. The effect has been explored across a broad range of material platforms and device geometries, including ferromagnet/normal-metal bilayers and low-dimensional systems [3, 7, 8]. This versatility has made spin pumping a widely used route for characterizing spin–charge conversion and interfacial spin transport [7, 9, 10].
@@
Recently, it has been shown that magnetization dynamics can induce strongly nonlinear transport responses, including the generation of ultrahigh harmonics in spin and charge pumping when additional symmetry-breaking mechanisms are present [11, 12]. Such high-harmonic spin-current emission naturally connects spin pumping to the broader field of ultrafast and nonlinear spintronics, where spin currents driven on femtosecond to picosecond timescales enable efficient terahertz radiation and frequency conversion [13, 14, 15, 16]. From this perspective, spin pumping emerges as a platform at the intersection of high-harmonic generation (HHG)[17] and spintronic terahertz emission [18], underscoring the intrinsically multidisciplinary nature of dynamical carrier transport phenomena.
In its conventional realization, spin pumping is driven by ferromagnetic [2] or antiferromagnetic [19] resonance, and the emitted spin current oscillates at the same frequency as the driving dynamics. Symmetry considerations prohibit direct charge transport in this setting, such that only pure spin angular momentum is pumped. In our previous works, we demonstrated that extremely high harmonics can emerge in carrier pumping in the presence of relativistic mechanisms [11] or noncollinear magnetic textures [20]. More recently, the altermagnetic order has been identified as a nonrelativistic platform capable of supporting high-harmonic carrier pumping [21], owing to its intrinsically nonrelativistic alternating spin–momentum coupling[22, 23].
@@
In this letter, we propose a distinct route to high-harmonic spin pumping that does not rely on spin-momentum coupling. Building on recent insights into the role of nonlinearities in adiabatic energy spectra [12], we show that the essential ingredient underlying high-harmonic pumping is not SOC per se, but rather the presence of intrinsic nonlinear band dynamics. While such nonlinearities often arise in systems with spin–orbit interactions, we demonstrate that they can also be generated in purely magnetic systems. Specifically, we consider a SOC-free magnetic structure in which a second static order parameter is introduced perpendicular to the cone axis of the precessing magnetization. This additional magnetic order qualitatively reshapes the adiabatic energy levels, inducing strong nonlinear features that directly propagate into the dynamical transport response. As a result, spin pumping accompanied by a cascade of higher harmonics becomes possible even in the complete absence of SOC. Our findings establish a new mechanism for nonlinear and high-frequency spin pumping rooted solely in magnetic structure and dynamics, thereby broadening the class of systems capable of hosting ultrafast spin-transport phenomena beyond the existing spin–orbit-based approaches.
II From Band Dynamics to Nonlinear Spin Transport
@@
In conventional spin pumping, a precessing magnetic order placed in proximity to a normal metal injects a spin current into the latter. From a band-structure perspective, this process is governed by the adiabatic evolution of electronic states in a time-dependent magnetic background. To identify the conditions under which this evolution can generate high-harmonic responses in transport, it is sufficient to consider a minimal model consisting of a one-dimensional chain hosting a driven magnetic order. Despite its simplicity, this model provides a transparent platform to connect nonlinear band dynamics to high-frequency spin transport.
@@
III Numerical Simulations
Figure 2: Schematic of the simulated geometry, comprising a precessing ferromagnetic order (blue cone) adjacent to a static magnetic order (red arrow) oriented perpendicular to the precession axis of the dynamical magnetization. The system is connected to a normal-metal lead (gray region), into which spin currents are pumped.
To complement the analysis of the energy spectrum presented above, we now turn to numerical simulations of the time-dependent transport response generated by the Hamiltonian in Eq. (3). In particular, we consider a one-dimensional chain hosting a driven magnetic order in addition to a perpendicular static magnetic order (see Fig. 2). We compute the resulting spin and charge currents flowing into an adjacent normal-metal lead (depicted in gray in Fig. 2). This approach allows us to directly access the full time-resolved transport response beyond analytical limits and to validate the emergence of nonlinear pumping signatures discussed previously.
@@ III Numerical Simulations
${\rm{I}}^{\sigma_{j}}(t)=\Im\!\left[\psi_{i}^{\dagger}(t)\,\mathcal{H}_{il}(t)\,\sigma_{j}\,\psi_{l}(t)\right],$ (5)
@@
$\boldsymbol{\rm{I}}^{\sigma}(t)=\frac{\hbar}{4\pi}\big(A_{r}\,\boldsymbol{m}(t)\times\dot{\boldsymbol{m}}(t)-A_{i}\,\dot{\boldsymbol{m}}(t)\big),$ (6)
@@
We now turn to the case where the static transverse magnetic field is included. Figure 3 shows the time-domain signals and corresponding Fourier spectra for two different precession cone angles. In agreement with spin-pumping theory, the dc charge current remains zero [see Figs. 3(b) and 3(d)]. However, the presence of the transverse field gives rise to a strongly nonlinear dynamical response, as evidenced by the emergence of higher harmonics in the Fourier spectra shown in Figs. 3(a) and 3(c).
Note that the dc spin current polarized parallel to the static field vanishes. In the absence of SOC, spin dynamics in the static-field region is governed by coherent precession about the field axis, which does not mix transverse spin components into the longitudinal one. As the pumped spin current injected from the dynamical region carries no dc polarization along the static-field direction, the corresponding spin component remains uncoupled and decays diffusively. Consequently, in the periodic steady state, the dc spin current parallel to the static field is strictly zero. In contrast, the spin currents polarized along the $y$ and $z$ directions acquire finite dc contributions. For a small precession cone angle, $\theta=\pi/8$, only a limited number of harmonics is observed. When the dynamical magnetization maximally couples to the secondary magnetic order, corresponding to $\theta=\pi/2$, the harmonic spectrum extends up to the 30th order [Fig. 3(c)].
@@
Figure 4: Calculated dc components of the pumped spin current polarized along the $y$ [(a)] and $z$ [(b)] directions are shown as functions of the driving frequency $\omega$ for precession cone angles ranging from $\theta=\pi/8$ to $\pi/2$. The pumped currents are evaluated for the parameters $J_{0}=\gamma$ and $J_{1}=2\gamma$. The dashed brown and black lines correspond to linear fits of the underlying data.
IV Discussion and Conclusion
@@
Another notable feature of this mechanism is the emergence of dc spin components in selected polarization channels, while the dc charge current vanishes. This behavior clearly distinguishes the present nonlinear magnetic pumping regime from conventional adiabatic spin pumping and provides an experimentally accessible signature of the underlying nonlinear dynamics.
Taken together, these results establish a conceptually distinct route to high-harmonic spin pumping driven purely by magnetic interactions. Although demonstrated here for uniform magnetic dynamics, the mechanism naturally extends to spatially extended magnetic textures in $d$ dimensions, where the interplay between temporal dynamics and spatial variations of the magnetic order introduces additional nonlinearities. These nonlinearities enable the redistribution of spectral weight from the fundamental drive frequency to higher harmonics, thereby enriching the spin-pumping response. More generally, the coexistence of multiple magnetic regions and time-dependent exchange fields suggests that nonlinear spin transport of this kind should be realizable in a broad class of magnetic systems, providing a promising platform for ultrafast spintronic functionalities and high-frequency signal generation.